The length of a new room that you are going to live in is 9.2 times bigger than its width. If the perimeter cannot be more than 115' what are the dimensions of the room?
I am lost and do not know what to do.

Answer by rothauserc(4718)

let w be the width of the room and l be the length, we are given that
l = 9.2 * w
the perimeter = 2*l + 2*w
115 = 2*(9.2*w) + 2*w
115 = 18.4w + 2w
20.4w = 115
w = 115 / 20.4 = 5.637254902 approx 5.64 feet
l = 9.2 * 5.64 = 51.888 approx 51.89 feet